OpenStudy (anonymous):

is the value of f(-3)=8 minimum or maximum

OpenStudy (anonymous):

what is f()...?

OpenStudy (anonymous):

f(-3)=8

OpenStudy (anonymous):

I mean... what is the function f =? not numeric, but as a function...

OpenStudy (anonymous):

other wise this is just define a point

OpenStudy (anonymous):

yeah but I got it anyway it was minimum but I would love to ask you a few more questions if I may

OpenStudy (anonymous):

shoot... thought Im confused how the point could min of itself in a space...? Ooo well. Ask away :)

OpenStudy (anonymous):

It was Max sorry and my other question was In Algebra the number of tickets sold each day for an upcoming performance of Handel’s Messiah is given by N(x)=-0.5x^2+10+10, where x is the number of days since the concert was first announced. When will daily ticket sales peak and how many tickets will be sold that day?

OpenStudy (anonymous):

did you mean: N(x)=-0.5x^2+10x+10 ?

OpenStudy (anonymous):

do you know differentiation? what it should be solved?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

N'(x)=-x+10 peak means N'(x)=0

OpenStudy (anonymous):

so, x=10

OpenStudy (anonymous):

Never mind that question I will ask some one else but do you how to Give exact and approximate solutions to three decimal places. x^2-5x+5=0 The exact solutions are x= the approximate solutions to 3 decimal places are x=

OpenStudy (anonymous):

exact: \[x=(5+/- \sqrt{5})/2\]

OpenStudy (anonymous):

approximate: x1=3.618 x2=1.382

OpenStudy (anonymous):

thanks any way but I am looking for one Approximate and you need to be more specific with your exact if I put the exact the way you did it would be wrong I don't know if it has to be sqrt 5 and -sqrt 5 over 2